OpenStudy (anonymous):

Solve 6x2 = 18x. x = 0, x = −3 x = 0, x = 3 x = 6, x = −3 x = 6, x = 3

OpenStudy (helder_edwin):

u can use factoring

OpenStudy (helder_edwin):

\[ \large 6x^2=18x \] \[ \large 6x^2-18x=0 \] \[ \large 6x(x-3)=0 \]

OpenStudy (anonymous):

simply divide both sides by x 6x=18 x=3 and ofcourse also 0

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

x=3 and x=0, use quadratic equation to solve setting C as "0"

OpenStudy (anonymous):

thank you too..! ^^ :)

OpenStudy (anonymous):

x=0,x=3

OpenStudy (anonymous):

thx :)